OpenStudy (anonymous):

3. Riley finds out that Cooper’s Storage and Shipping Company is working with a local business to package some office supplies. Some of the supplies are packed inside a cube-shaped box with side lengths of 4 1/2 in. These boxes are then packed into a shipping box with dimensions of 18 in. × 9 in. × 4 1/2 in. How many boxes of office supplies can be packed into the larger box for shipping? Show your work.

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

First find the volume of a cube-shaped box then find the volume of the box they go into. Remember the sides of a cube are all the same length. 4.5*4.5*4.5 = 21.125 18*9*4.5 = 729 Now divide the volume of the larger box by the volume of the smaller box to see how many can go in it. :) 729 / 21.125 = 34 Your answer is 34!

OpenStudy (anonymous):

thanks one more:)

OpenStudy (anonymous):

4. One of the offices at the company was being repainted. Riley’s father challenged him to figure out how many quarts of paint would be needed. The dimensions of the office are 16 ft by 12 ft by 10 ft. The office floor, ceiling, and door would not be painted and there are no windows in this office. The door opening measures 3 ft by 8 ft. One quart of paint covers 100 ft2. How many quarts of paint would be needed to paint the room, if only full quarts can be purchased?

OpenStudy (kamibug):

I think the total area of the four walls is 700ft^2. But, I'm not sure b/c the question doesn't specify which of the dimensions are height, width, or length... The area of the door is 3*8=24 Subtract that from the total wall area: 700-24=676 Divide that by 100 to find how many quarts you'll need. 676 / 100 = 7 quarts But, again, this might be wrong because I'm not sure I calculated the total wall area correctly, sorry. :[

OpenStudy (anonymous):

its okay:)